can someone help me solve: given logb3 = 0.6826 
and logb7 = 1.2091, evaluate logb21. 
completely lost.

Answer by Edwin McCravy(20060)


can someone help me solve: given logb3 = 0.6826 
and logb7 = 1.2091, evaluate logb21. 
completely lost.

You need to know this basic rule of logarithms:

logb(X·Y) = logbX + logbY

and the fact that 21 = 3·7, to get

logb(3·7) = logb3 + logb7

then since you are given that 

logb3 = 0.6826 and logb7 = 1.2091

logb(3·7) = logb3 + logb7 = 0.6826 + 1.2091 = 1.8917

Edwin